OpenStudy (anonymous):

factor the quadtric expression completely and find the roots for the expression 20x+13x+2

OpenStudy (anonymous):

u mean 20x^2?

OpenStudy (anonymous):

( 5x + 2 )( 4x + 1 ) the roots if this = 0 are -2/5 , -1/4

OpenStudy (anonymous):

i assumed u meant 20x^2

OpenStudy (anonymous):

it is not a quad equation!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you find those roots @jimmy rep

OpenStudy (anonymous):

first factorised to get 2 factors when two items multiplied together = 0 then one of them must also = 0

OpenStudy (amistre64):

20x^2+13x+2 20(2) = 40 1,40 2,20 4,10 8,5 ... 8 +5 = 13 .... (x+8/20) (x+5/20) ; reduce (x+2/5) (x+1/4) ; stick bottoms in front (5x+2) (4x+1)

OpenStudy (amistre64):

to determine the roots we can stop at: (x+2/5) (x+1/4) ; and solve for 0 :) x = -2/5 and x = -1/4